[figure: 9.7 cm]

MAGNETIC ANISOTROPY IN EuS-PbS MULTILAYERS

T. Story^{1}, C.H.W. Swüste^{2}, H.J.M. Swagten^{2}, W.J.M. de Jonge^{2} A. Stachow-Wójcik^{3}, A. Twardowski^{3}, M. Arciszewska^{1}, W. Dobrowolski^{1}, R.R. Gała̧zka^{1}, A.Yu. Sipatov^{4} 1Institute of Physics Polish Academy of Sciences, Warsaw, Poland 2Eindhoven University of Technology, Eindhoven, The Netherlands 3Institute of Experimental Physics, Warsaw University, Warsaw, Poland 4Kharkov State Polytechnical University, Kharkov, Ukraine

The EuS-PbS multilayers form semiconducting ferromagnet-diamagnet structures which are model low dimensional Heisenberg ferromagnets. The ferromagnetic transition is observed in the structures with EuS thickness d down to 2 monolayers (ML). For EuS thicknesses smaller than about 10ML, the Curie temperature decreases with decreasing EuS thickness, ΔT_{c} ∼ −1/d. The transition temperature also depends on stress present in the structures. We present the results of ferromagnetic resonance studies of the thickness dependence of magnetic anisotropy in 2 series of EuS-PbS multilayers grown on (111) BaF_{2} and (100) KCl substrates with the EuS thickness varying in the range d = 2 − 50ML. The anisotropy constant K was found to follow the dependence K(d) = K_{v} + K_{s}/d, with the surface term K_{s} larger for layers grown on BaF_{2} as compared to KCl. This difference is discussed in terms of different thermal stress-induced distortions of cubic crystal lattice of EuS. We found that the thickness of EuS layer required for the perpendicular (to the layer) magnetization is d ≤ 2 Å, i.e., it is below 1ML.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : oral

Corresponding author : Tomasz Story

Address for correspondence : Instytut Fizyki PAN Al. Lotników 32/46 02-668 Warszawa

Email address : story@ifpan.edu.pl